OpenStudy (anonymous):

Simplify: ln e^x + 4y

OpenStudy (jit4won):

check your question

OpenStudy (anonymous):

its typed correctly

OpenStudy (experimentx):

\( \ln e^a = a\)

OpenStudy (anonymous):

a. x/4y b.x + 4y c.e^x + 4y d.4xy

OpenStudy (anonymous):

possible answers ^

OpenStudy (experimentx):

use the above property on \( \ln e^x\)

OpenStudy (jit4won):

b x+4Y

OpenStudy (anonymous):

x+4y